OpenStudy (anonymous):

please solve -x/3 = -3/4

OpenStudy (anonymous):

the minuses cancel out

OpenStudy (lgbasallote):

\[-\frac x 3 = -\frac 34\] just multiply both sides by -3 \[x = -\frac 34 \times -3\]

OpenStudy (anonymous):

x -3 /1?

OpenStudy (lgbasallote):

\[-\frac 34 \times -3 \implies \frac{-3 \times -3}{4 \times 1}\]

OpenStudy (anonymous):

=9/4?

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

thanks!